# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ Appendix D Technical Lemmas
$\hat{\mathbf{U}}\mathbf{W}-\mathbf{U}$ $=\mathbf{E}\mathbf{U}\bm{\Lambda}^{-1}+\underbrace{\mathbf{U}(\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top})\mathbf{W}}_{\mathbf{R}_{\mathbf{U},1}}+\underbrace{\mathbf{U}(\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}})\hat{\bm{\Lambda}}^{-1}\mathbf{W}}_{\mathbf{R}_{\mathbf{U},2}}$
$+\underbrace{\mathbf{E}\mathbf{U}\bm{\Lambda}^{-1}(\bm{\Lambda}\mathbf{W}^{\top}-\mathbf{W}^{\top}\hat{\bm{\Lambda}})\hat{\bm{\Lambda}}^{-1}\mathbf{W}}_{\mathbf{R}_{\mathbf{U},3}}+\underbrace{\mathbf{E}(\hat{\mathbf{U}}\mathbf{W}-\mathbf{U})\mathbf{W}^{\top}\hat{\bm{\Lambda}}^{-1}\mathbf{W}}_{\mathbf{R}_{\mathbf{U},4}}.$
Let $\mathbf{W}_{\mathbf{U}}$ be the minimizer of $\|\hat{\mathbf{U}}\mathbf{O}-\mathbf{U}\|_{F}$ over all $d\times d$ orthogonal matrices $\mathbf{O}$. We now bound the spectral norms of $\mathbf{R}_{\mathbf{U},1},\dots,\mathbf{R}_{\mathbf{U},4}$ when $\mathbf{W}=\mathbf{W}_{\mathbf{U}}$.
For $\mathbf{R}_{\mathbf{U},1}$, by Lemma D.2 we have
$\|\mathbf{R}_{\mathbf{U},1}\|\leq\|\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$
$\|\mathbf{R}_{\mathbf{U},1}\|_{2\to\infty}\leq\|\mathbf{U}\|_{2\to\infty}\cdot\|\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}}{q_{i}\lambda_{i,\min}^{2}}$
@@
For $\mathbf{R}_{\mathbf{U},2}$, by Lemma D.2 and Lemma D.3 we have
$\|\mathbf{R}_{\mathbf{U},2}\|\leq$ $\|\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}^{(i)}\|\cdot\|\hat{\bm{\Lambda}}^{-1}\|$
@@
$\|\mathbf{R}_{\mathbf{U},2}\|_{2\to\infty}$ $\leq\|\mathbf{U}\|_{2\to\infty}\cdot\|\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}^{(i)}\|\cdot\|\hat{\bm{\Lambda}}^{-1}\|$
@@
For $\mathbf{R}_{\mathbf{U},3}$, by Lemma D.1, Lemma D.2 and Lemma D.3 we have
$\|\mathbf{R}_{\mathbf{U},3}\|$ $\leq\|\mathbf{E}\|\cdot\|\bm{\Lambda}^{-1}\|\cdot\|\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}^{\top}_{\mathbf{U}}\hat{\bm{\Lambda}}\|\cdot\|\hat{\bm{\Lambda}}^{-1}\|$
@@
$\|\mathbf{R}_{\mathbf{U},3}\|_{2\to\infty}$ $\leq\|\mathbf{E}\mathbf{U}\|_{2\to\infty}\cdot\|\bm{\Lambda}^{-1}\|\cdot\|\bm{\Lambda}\mathbf{W}^{\top}_{\mathbf{U}}-\mathbf{W}^{\top}_{\mathbf{U}}\hat{\bm{\Lambda}}\|\cdot\|\hat{\bm{\Lambda}}^{-1}\|$
@@
For $\mathbf{R}_{\mathbf{U},4}$, by Lemma D.1, Lemma D.2 and Lemma D.9, we have
$\|\mathbf{R}_{\mathbf{U},4}\|$ $\leq\|\mathbf{E}\|\cdot\|\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U}\|\cdot\|\hat{\bm{\Lambda}}^{-1}\|$
@@
$\|\mathbf{R}_{\mathbf{U},4}\|_{2\to\infty}$ $\leq\|\mathbf{E}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|_{2\to\infty}\cdot\|\hat{\bm{\Lambda}}^{-1}\|$
@@
with high probability as $\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}\lesssim 1$ as implied by Eq. (C.3).
D.4 Technical lemmas for $\mathbf{R}_{\mathbf{U},4}$ in Lemma D.4
Our bound for $\mathbf{R}_{\mathbf{U},4}$ in the above proof of Lemma D.4 is based on a series of technical lemmas which culminate in a high-probability bound for $\|\mathbf{E}^{(i)}(\hat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)})\|_{2\to\infty}$. These lemmas are derived using an adaptation of the leave-one-out analysis presented in Theorem 3.2 of Xie (2024); the noise model for $\mathbf{N}^{(i)}$ in the current paper is, however, somewhat different from that of Xie (2024) and thus we chose to provide self-contained proofs of these lemmas here. Once again, for ease of exposition we will fix a value of $i$ and thereby drop this index from our matrices in this section.
@@
$\mathbf{A}_{s,t}^{[h]}=\begin{cases}\mathbf{A}_{s,t},&\text{ if }s\neq h\text{and }t\neq h,\\ \mathbf{P}_{s,t},&\text{ if }s=h\text{ or }s=h.\end{cases}$
@@
$\mathbf{A}=\hat{\mathbf{U}}\hat{\bm{\Lambda}}\hat{\mathbf{U}}^{\top}+\hat{\mathbf{U}}_{\perp}\hat{\bm{\Lambda}}_{\perp}\hat{\mathbf{U}}^{\top}_{\perp},$
$\mathbf{A}^{[h]}$ $=\hat{\mathbf{U}}^{[h]}\hat{\bm{\Lambda}}^{[h]}\hat{\mathbf{U}}^{[h]\top}+\hat{\mathbf{U}}^{[h]}_{\perp}\hat{\bm{\Lambda}}^{[h]}_{\perp}\hat{\mathbf{U}}^{[h]\top}_{\perp}.$
Lemma D.5.
Consider the setting in Lemma C.1, we then have
$\|\hat{\mathbf{U}}^{(i)}\|_{2\to\infty}\lesssim\frac{d^{1/2}}{n_{i}^{1/2}},\quad\|(\hat{\mathbf{U}}^{(i)})^{[h]}\|_{2\to\infty}\lesssim\frac{d^{1/2}}{n_{i}^{1/2}}$
with high probability. Furthermore, let $\mathbf{W}^{[h]}$ be the solution of orthogonal Procrustes problem between $\hat{\mathbf{U}}^{[h]}$ and $\mathbf{U}$. We then have
$\|(\hat{\mathbf{U}}^{(i)})^{[h]}(\mathbf{W}^{(i)})^{[h]}-\mathbf{U}^{(i)}\|_{2\to\infty}\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
The proof is based on verifying the conditions in Theorem 2.1 of Abbe et al. (2020), and this can be done by following the exact same derivations as that in Lemma 12 of Abbe et al. (2020). More specifically, $\Delta^{*}$ in Abbe et al. (2020) corresponds to $\lambda_{i,\min}$ in our paper while $\kappa$ in Abbe et al. (2020) corresponds to $M$ in our setting, where $M$ appears in the assumptions of Lemma C.1 and is bounded, and $\gamma$ in Abbe et al. (2020) can be set to be $\frac{c(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$ for some sufficiently large constant $c>0$, based on the bound of $\|\mathbf{E}\|$ from Lemma D.1. Then all desired results of Lemma D.5 can by obtained under the conditions that $q_{i}n_{i}\gg\log n_{i}$ and condition in Eq. (C.3).
Lemma D.6.
Consider the setting in Lemma C.1. Recall in Lemma D.1, we decompose $\mathbf{E}^{(i)}$ as $\mathbf{E}^{(i,1)}+\mathbf{E}^{(i,2)}$. Let $\mathbf{e}^{(i)}_{h}$ denote the $h$th row of $\mathbf{E}^{(i)}$. Then for any deterministic $n_{i}\times d$ matrix $\mathbf{C}$, we have
@@
For any $h\in[n_{i}]$, let $\mathbf{e}^{(i,1)}_{h}$ and $\mathbf{e}^{(i,2)}_{h}$ denote the $h$th row of $\mathbf{E}^{(i,1)}$ and $\mathbf{E}^{(i,2)}$. Following the same arguments as that for $\|\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}\|_{2\to\infty}$ and $\|\mathbf{E}^{(i,2)}\mathbf{U}^{(i)}\|_{2\to\infty}$ in the proof of Lemma D.1, we have
@@
Lemma D.7.
Consider the setting in Lemma D.4, we then have
$\|\sin\Theta((\hat{\mathbf{U}}^{(i)})^{[h]},\hat{\mathbf{U}}^{(i)})\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
By the construction of $\mathbf{A}^{[h]}$ and Lemma D.1 we have
@@
$\|\mathbf{A}^{[h]}-\mathbf{P}\|\leq\|\mathbf{A}^{[h]}-\mathbf{A}\|+\|\mathbf{E}\|\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ (D.15)
@@
$|\lambda_{d}(\mathbf{A}^{[h]})-\lambda_{i,\min}|\leq\|\mathbf{A}^{[h]}-\mathbf{P}\|\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ (D.16)
with high probability. The condition in Eq. (C.3) implies $\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}=o(1)$ and hence $\lambda_{d}(\mathbf{A}^{[h]})\asymp\lambda_{i,\min}$. Furthermore, by Lemma D.2 we have $\lambda_{d+1}(\mathbf{A})\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ with high probability. Applying Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 of Stewart and Sun (1990)) we have
$\|\sin\Theta(\hat{\mathbf{U}}^{[h]},\hat{\mathbf{U}})\|\leq\frac{\|(\mathbf{A}^{[h]}-\mathbf{A})\hat{\mathbf{U}}^{[h]}\|}{\lambda_{d}(\mathbf{A}^{[h]})-\lambda_{d+1}(\mathbf{A})}\lesssim\frac{\|(\mathbf{A}^{[h]}-\mathbf{A})\hat{\mathbf{U}}^{[h]}\|_{F}}{\lambda_{i,\min}}$ (D.17)
with high probability. We now bound $\|(\mathbf{A}^{[h]}-\mathbf{A})\hat{\mathbf{U}}^{[h]}\|_{F}$. Note that
$\|(\mathbf{A}^{[h]}-\mathbf{A})\hat{\mathbf{U}}^{[h]}\|_{F}$ $=\Big{[}\sum_{s\in[n_{i}],s\neq h}\sum_{r\in[d]}(\mathbf{E}_{s,h}\hat{\mathbf{U}}^{[h]}_{h,r})^{2}+\Big{\|}\mathbf{e}_{h}^{\top}\hat{\mathbf{U}}^{[h]}\Big{\|}^{2}\Big{]}^{1/2},$ (D.18)
where $\mathbf{e}_{h}$ is the $h$th row of $\mathbf{E}$. For the first term on the right side of Eq. (D.18), by Cauchy-Schwarz inequality, Lemma D.5 and Lemma D.1 we have
$\big{[}\sum_{s\in[n_{i}],s\neq h}\sum_{r\in[d]}(\mathbf{E}_{s,h}\hat{\mathbf{U}}^{[h]}_{h,r})^{2}\big{]}^{1/2}$ $\leq\|\mathbf{E}\|_{2\to\infty}\cdot\|\hat{\mathbf{U}}^{[h]}\|_{2\to\infty}$ (D.19)
$\leq\|\mathbf{E}\|\cdot\|\hat{\mathbf{U}}^{[h]}\|_{2\to\infty}\lesssim d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})$
with high probability. For the second term on the right side of Eq. (D.18), as $\mathbf{e}_{h}$ and $\hat{\mathbf{U}}^{[h]}$ are independent, we have by Lemma D.6, Lemma D.5, and the assumption $n_{i}q_{i}\gtrsim\log^{2}n_{i}$ that
$\|\mathbf{e}_{h}^{\top}\hat{\mathbf{U}}^{[h]}\|$ $\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}\|\hat{\mathbf{U}}^{[h]}\|_{2\to\infty}\log^{1/2}n_{i}+q_{i}^{-1}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i}\log^{1/2}n_{i})\|\hat{\mathbf{U}}^{[h]}\|_{2\to\infty}\log n_{i}$ (D.20)
@@
with high probability. Combining Eq. (D.18), Eq. (D.19) and Eq. (D.20), we have
$\|(\mathbf{A}^{[h]}-\mathbf{A})\hat{\mathbf{U}}^{[h]}\|_{F}\lesssim d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}$ (D.21)
with high probability. Substituting Eq. (D.21) into Eq. (D.17) yields the desired claim. ∎
Lemma D.8.
Consider the setting in Lemma D.4, we then have
$\|(\mathbf{e}_{h}^{(i)})^{\top}[(\hat{\mathbf{U}}^{(i)})^{[h]}(\hat{\mathbf{U}}^{(i)})^{[h]\top}\mathbf{U}^{(i)}-\mathbf{U}^{(i)}]\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$

### v4

@@ Appendix E Technical Lemmas
$\widehat{\mathbf{U}}\mathbf{W}-\mathbf{U}$ $=\mathbf{E}\mathbf{U}\bm{\Lambda}^{-1}+\underbrace{\mathbf{U}(\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{W}^{\top})\mathbf{W}}_{\mathbf{R}_{\mathbf{U},1}}+\underbrace{\mathbf{U}(\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}})\widehat{\bm{\Lambda}}^{-1}\mathbf{W}}_{\mathbf{R}_{\mathbf{U},2}}$
$+\underbrace{\mathbf{E}\mathbf{U}\bm{\Lambda}^{-1}(\bm{\Lambda}\mathbf{W}^{\top}-\mathbf{W}^{\top}\widehat{\bm{\Lambda}})\widehat{\bm{\Lambda}}^{-1}\mathbf{W}}_{\mathbf{R}_{\mathbf{U},3}}+\underbrace{\mathbf{E}(\widehat{\mathbf{U}}\mathbf{W}-\mathbf{U})\mathbf{W}^{\top}\widehat{\bm{\Lambda}}^{-1}\mathbf{W}}_{\mathbf{R}_{\mathbf{U},4}}.$
Let $\mathbf{W}_{\mathbf{U}}$ be the minimizer of $\|\widehat{\mathbf{U}}\mathbf{O}-\mathbf{U}\|_{F}$ over all $d\times d$ orthogonal matrices $\mathbf{O}$. We now bound the spectral norms of $\mathbf{R}_{\mathbf{U},1},\dots,\mathbf{R}_{\mathbf{U},4}$ when $\mathbf{W}=\mathbf{W}_{\mathbf{U}}$.
For $\mathbf{R}_{\mathbf{U},1}$, by Lemma E.2 we have
$\|\mathbf{R}_{\mathbf{U},1}\|\leq\|\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{W}^{\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$
$\|\mathbf{R}_{\mathbf{U},1}\|_{2\to\infty}\leq\|\mathbf{U}\|_{2\to\infty}\cdot\|\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{W}^{\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}}{q_{i}\lambda_{i,\min}^{2}}$
@@
For $\mathbf{R}_{\mathbf{U},2}$, by Lemma E.2 and Lemma E.3 we have
$\|\mathbf{R}_{\mathbf{U},2}\|\leq$ $\|\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}^{(i)}\|\cdot\|\widehat{\bm{\Lambda}}^{-1}\|$
@@
$\|\mathbf{R}_{\mathbf{U},2}\|_{2\to\infty}$ $\leq\|\mathbf{U}\|_{2\to\infty}\cdot\|\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}^{(i)}\|\cdot\|\widehat{\bm{\Lambda}}^{-1}\|$
@@
For $\mathbf{R}_{\mathbf{U},3}$, by Lemma E.1, Lemma E.2 and Lemma E.3 we have
$\|\mathbf{R}_{\mathbf{U},3}\|$ $\leq\|\mathbf{E}\|\cdot\|\bm{\Lambda}^{-1}\|\cdot\|\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}^{\top}_{\mathbf{U}}\widehat{\bm{\Lambda}}\|\cdot\|\widehat{\bm{\Lambda}}^{-1}\|$
@@
$\|\mathbf{R}_{\mathbf{U},3}\|_{2\to\infty}$ $\leq\|\mathbf{E}\mathbf{U}\|_{2\to\infty}\cdot\|\bm{\Lambda}^{-1}\|\cdot\|\bm{\Lambda}\mathbf{W}^{\top}_{\mathbf{U}}-\mathbf{W}^{\top}_{\mathbf{U}}\widehat{\bm{\Lambda}}\|\cdot\|\widehat{\bm{\Lambda}}^{-1}\|$
@@
For $\mathbf{R}_{\mathbf{U},4}$, by Lemma E.1, Lemma E.2 and Lemma E.9, we have
$\|\mathbf{R}_{\mathbf{U},4}\|$ $\leq\|\mathbf{E}\|\cdot\|\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U}\|\cdot\|\widehat{\bm{\Lambda}}^{-1}\|$
@@
$\|\mathbf{R}_{\mathbf{U},4}\|_{2\to\infty}$ $\leq\|\mathbf{E}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\|_{2\to\infty}\cdot\|\widehat{\bm{\Lambda}}^{-1}\|$
@@
with high probability as $\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}\lesssim 1$ as implied by Eq. (D.3).
E.4 Technical lemmas for $\mathbf{R}_{\mathbf{U},4}$ in Lemma E.4
Our bound for $\mathbf{R}_{\mathbf{U},4}$ in the above proof of Lemma E.4 is based on a series of technical lemmas which culminate in a high-probability bound for $\|\mathbf{E}^{(i)}(\widehat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)})\|_{2\to\infty}$. These lemmas are derived using an adaptation of the leave-one-out analysis presented in Theorem 3.2 of Xie (2024); the noise model for $\mathbf{N}^{(i)}$ in the current paper is, however, somewhat different from that of Xie (2024) and thus we chose to provide self-contained proofs of these lemmas here. Once again, for ease of exposition we will fix a value of $i$ and thereby drop this index from our matrices in this section.
@@
[…]
@@
$\mathbf{A}=\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}\widehat{\mathbf{U}}^{\top}+\widehat{\mathbf{U}}_{\perp}\widehat{\bm{\Lambda}}_{\perp}\widehat{\mathbf{U}}^{\top}_{\perp},$
$\mathbf{A}^{[h]}$ $=\widehat{\mathbf{U}}^{[h]}\widehat{\bm{\Lambda}}^{[h]}\widehat{\mathbf{U}}^{[h]\top}+\widehat{\mathbf{U}}^{[h]}_{\perp}\widehat{\bm{\Lambda}}^{[h]}_{\perp}\widehat{\mathbf{U}}^{[h]\top}_{\perp}.$
Lemma E.5.
Consider the setting in Lemma D.1, we then have
$\|\widehat{\mathbf{U}}^{(i)}\|_{2\to\infty}\lesssim\frac{d^{1/2}}{n_{i}^{1/2}},\quad\|(\widehat{\mathbf{U}}^{(i)})^{[h]}\|_{2\to\infty}\lesssim\frac{d^{1/2}}{n_{i}^{1/2}}$
with high probability. Furthermore, let $\mathbf{W}^{[h]}$ be the solution of orthogonal Procrustes problem between $\widehat{\mathbf{U}}^{[h]}$ and $\mathbf{U}$. We then have
$\|(\widehat{\mathbf{U}}^{(i)})^{[h]}(\mathbf{W}^{(i)})^{[h]}-\mathbf{U}^{(i)}\|_{2\to\infty}\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
The proof is based on verifying the conditions in Theorem 2.1 of Abbe et al. (2020), and this can be done by following the exact same derivations as that in Lemma 12 of Abbe et al. (2020). More specifically, $\Delta^{*}$ in Abbe et al. (2020) corresponds to $\lambda_{i,\min}$ in our paper while $\kappa$ in Abbe et al. (2020) corresponds to $M$ in our setting, where $M$ appears in the assumptions of Lemma D.1 and is bounded, and $\gamma$ in Abbe et al. (2020) can be set to be $\frac{c(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$ for some sufficiently large constant $c>0$, based on the bound of $\|\mathbf{E}\|$ from Lemma E.1. Then all desired results of Lemma E.5 can by obtained under the conditions that $q_{i}n_{i}\gg\log n_{i}$ and condition in Eq. (D.3).
Lemma E.6.
Consider the setting in Lemma D.1. Recall in Lemma E.1, we decompose $\mathbf{E}^{(i)}$ as $\mathbf{E}^{(i,1)}+\mathbf{E}^{(i,2)}$. Let $\mathbf{e}^{(i)}_{h}$ denote the $h$th row of $\mathbf{E}^{(i)}$. Then for any deterministic $n_{i}\times d$ matrix $\mathbf{C}$, we have
@@
For any $h\in[n_{i}]$, let $\mathbf{e}^{(i,1)}_{h}$ and $\mathbf{e}^{(i,2)}_{h}$ denote the $h$th row of $\mathbf{E}^{(i,1)}$ and $\mathbf{E}^{(i,2)}$. Following the same arguments as that for $\|\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}\|_{2\to\infty}$ and $\|\mathbf{E}^{(i,2)}\mathbf{U}^{(i)}\|_{2\to\infty}$ in the proof of Lemma E.1, we have
@@
Lemma E.7.
Consider the setting in Lemma E.4, we then have
$\|\sin\Theta((\widehat{\mathbf{U}}^{(i)})^{[h]},\widehat{\mathbf{U}}^{(i)})\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
By the construction of $\mathbf{A}^{[h]}$ and Lemma E.1 we have
@@
$\|\mathbf{A}^{[h]}-\mathbf{P}\|\leq\|\mathbf{A}^{[h]}-\mathbf{A}\|+\|\mathbf{E}\|\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ (E.15)
@@
$|\lambda_{d}(\mathbf{A}^{[h]})-\lambda_{i,\min}|\leq\|\mathbf{A}^{[h]}-\mathbf{P}\|\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ (E.16)
with high probability. The condition in Eq. (D.3) implies $\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}=o(1)$ and hence $\lambda_{d}(\mathbf{A}^{[h]})\asymp\lambda_{i,\min}$. Furthermore, by Lemma E.2 we have $\lambda_{d+1}(\mathbf{A})\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}$ with high probability. Applying Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 of Stewart and Sun (1990)) we have
$\|\sin\Theta(\widehat{\mathbf{U}}^{[h]},\widehat{\mathbf{U}})\|\leq\frac{\|(\mathbf{A}^{[h]}-\mathbf{A})\widehat{\mathbf{U}}^{[h]}\|}{\lambda_{d}(\mathbf{A}^{[h]})-\lambda_{d+1}(\mathbf{A})}\lesssim\frac{\|(\mathbf{A}^{[h]}-\mathbf{A})\widehat{\mathbf{U}}^{[h]}\|_{F}}{\lambda_{i,\min}}$ (E.17)
with high probability. We now bound $\|(\mathbf{A}^{[h]}-\mathbf{A})\widehat{\mathbf{U}}^{[h]}\|_{F}$. Note that
$\|(\mathbf{A}^{[h]}-\mathbf{A})\widehat{\mathbf{U}}^{[h]}\|_{F}$ $=\Big{[}\sum_{s\in[n_{i}],s\neq h}\sum_{r\in[d]}(\mathbf{E}_{s,h}\widehat{\mathbf{U}}^{[h]}_{h,r})^{2}+\Big{\|}\mathbf{e}_{h}^{\top}\widehat{\mathbf{U}}^{[h]}\Big{\|}^{2}\Big{]}^{1/2},$ (E.18)
where $\mathbf{e}_{h}$ is the $h$th row of $\mathbf{E}$. For the first term on the right side of Eq. (E.18), by Cauchy-Schwarz inequality, Lemma E.5 and Lemma E.1 we have
$\big{[}\sum_{s\in[n_{i}],s\neq h}\sum_{r\in[d]}(\mathbf{E}_{s,h}\widehat{\mathbf{U}}^{[h]}_{h,r})^{2}\big{]}^{1/2}$ $\leq\|\mathbf{E}\|_{2\to\infty}\cdot\|\widehat{\mathbf{U}}^{[h]}\|_{2\to\infty}$ (E.19)
$\leq\|\mathbf{E}\|\cdot\|\widehat{\mathbf{U}}^{[h]}\|_{2\to\infty}\lesssim d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})$
with high probability. For the second term on the right side of Eq. (E.18), as $\mathbf{e}_{h}$ and $\widehat{\mathbf{U}}^{[h]}$ are independent, we have by Lemma E.6, Lemma E.5, and the assumption $n_{i}q_{i}\gtrsim\log^{2}n_{i}$ that
$\|\mathbf{e}_{h}^{\top}\widehat{\mathbf{U}}^{[h]}\|$ $\lesssim q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}\|\widehat{\mathbf{U}}^{[h]}\|_{2\to\infty}\log^{1/2}n_{i}+q_{i}^{-1}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i}\log^{1/2}n_{i})\|\widehat{\mathbf{U}}^{[h]}\|_{2\to\infty}\log n_{i}$ (E.20)
@@
with high probability. Combining Eq. (E.18), Eq. (E.19) and Eq. (E.20), we have
$\|(\mathbf{A}^{[h]}-\mathbf{A})\widehat{\mathbf{U}}^{[h]}\|_{F}\lesssim d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}$ (E.21)
with high probability. Substituting Eq. (E.21) into Eq. (E.17) yields the desired claim. ∎
Lemma E.8.
Consider the setting in Lemma E.4, we then have
$\|(\mathbf{e}_{h}^{(i)})^{\top}[(\widehat{\mathbf{U}}^{(i)})^{[h]}(\widehat{\mathbf{U}}^{(i)})^{[h]\top}\mathbf{U}^{(i)}-\mathbf{U}^{(i)}]\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}^{1/2}\log n_{i}}{q_{i}\lambda_{i,\min}}$